# IndirectRequests: Making Task-Oriented Dialogue Datasets More Natural by Synthetically Generating Indirect User Requests

### v1

@@ Abstract
Existing benchmark corpora of task-oriented dialogue are collected either using a “machines talking to machines” approach or by giving template-based goal descriptions to crowdworkers. These methods, however, often produce utterances that are markedly different from natural human conversations in which people often convey their preferences in indirect ways, such as through small talk. We term such utterances as Indirect User Requests (IURs). Understanding such utterances demands considerable world knowledge and reasoning capabilities on the listener’s part. Our study introduces a large language model (LLM)-based pipeline to automatically generate realistic, high-quality IURs for a given domain, with the ultimate goal of supporting research in Natural Language Understanding (NLU) and Dialogue State Tracking (DST) for task-oriented dialogue systems. Our findings show that while large LLMs such as GPT-3.5 and GPT-4 generate high-quality IURs, achieving similar quality with smaller models is more challenging. We release IndirectRequests, a dataset of IURs that advances beyond the initial Schema Guided Dialog (SGD) dataset in that it provides a challenging testbed for testing the “in the wild” performance of NLU and DST models.
@@ 1 Introduction
Task-oriented dialogue assistants (Balaraman et al., 2021) help people carry out tasks such as making hotel reservations, setting alarms, looking up train schedules, and so on through natural language conversations (Budzianowski et al., 2018; Mosig et al., 2020; Byrne et al., 2019; Asri et al., 2017). One of the most challenging aspects of developing a dialogue assistant is developing the natural language understanding model (Mehri et al., 2020). With the proliferation of powerful LLMs (Brown et al., 2020), great strides have been made in model performance on a handful of benchmark datasets (Budzianowski et al., 2018; Rastogi et al., 2020; Asri et al., 2017) that are widely regarded as proxies for the general task. However, despite these advances, many models suffer from errors when presented with utterances that differ in particular ways from those present in their training datasets (Cho et al., 2021, 2022). As a result, users end up feeling frustrated when using talking to these virtual assistants freely as they would with other humans (Mavrina et al., 2022).
@@
There are several failure modes of NLU and DST models, one of them being the lack of model capability to understand indirect requests that do not directly mention the target slot value as expected by the system (Cohen, 2019). For example, while reserving a hotel room, rather than saying the presumably more natural utterance, “it’s gonna be me, my wife, and our twins”, [^1] This is presumably a natural thing to say during a friendly chat with a human receptionist. a user might instead resort to more direct terms (for example, “I want to book a room for four people”) to ensure that the intent of the utterance is understood by the virtual assistant on the first attempt. Figure 1 shows two notional instances of IURs in the context of a restaurant reservation and an intelligent home-assistant dialogue respectively.
From a machine learning standpoint, the challenge DST models encounter in understanding IURs stems from the lack of labeled examples in mainstream benchmark datasets used for developing task-oriented dialogue agents (Cho et al., 2021). Furthermore, the main reason for this discrepancy in distribution between benchmark datasets and “in-the-wild” utterances can be attributed to the controlled environment of laboratory settings where datasets are crowdsourced Zarcone et al. (2021).
Existing benchmark datasets of task-oriented dialogue such as MultiWOZ (Budzianowski et al., 2018), Schema Guided Dialog (SGD) (Rastogi et al., 2020), and FRAMES (Asri et al., 2017) all suffer from this distributional mismatch. For example, we found that only around 500 out of over 10,000 user utterances in the SGD dataset do not contain an explicit mention of the target slot value.
To bridge this distributional gap, we present an LLM-based data generation pipeline to scalably generate IURs for a new task-oriented dialogue domain. Our work contributes the following:
@@
    We suggest linguistic criteria to formalize the concept of what constitutes an indirect IUR. Specifically, we develop a set of linguistic criteria to systematically evaluate questions such as “What counts as an indirect user request?” and “How indirectly is this user request phrased with respect to a given domain schema?” in task-oriented dialogue contexts.
@@
    Finally, we empirically demonstrate the increased difficulty of the IURs by showing that the performance of a state-of-the-art DST model significantly degrades when applied on IndirectRequests utterances as compared to their counterparts from SGD.
@@
Before outlining the linguistic criteria, however, we first describe the paradigm of “schema-guided dialogue” since it serves as the basis for the criteria.
2 Schema-Guided Dialogue
@@ 4.3 Dataset Splits
Based on the crowdsourced labels for both Unambiguity and World-Understanding, we curate the IndirectRequests dataset and release it for public use. [^2] https://huggingface.co/datasets/msamogh/indirect-requests/ In going from the “raw” crowdsourced samples to the dataset, we split the dataset and systematically create labels for each sample for both Unambiguity and World-Understanding criteria. While splitting IndirectRequests into train, validation, and test sets, we split our samples based on same lines on which the services are split across the SGD dataset. This alignment with the SGD dataset splits is intended to aid future work that might need to compare our results with previous work reporting on SGD.
@@ 5.1 Proxy Evaluation Results
For the small LM category, we employ BERT-based models in a zero-shot setup. For the Unambiguity criterion, we frame the evaluation as $k$ Natural Language Inference (NLI) problems, where $k$ is the number of possible slot values. Each problem considers the candidate IUR as the premise and a possible slot value as the hypothesis. We use a BERT-based NLI model333nli-deberta-v3-small to obtain entailment scores and return the argmax score. If the maximum score is below 0.3, we deem the IUR ambiguous for that slot. For World-Understanding, we use ms-marco-MiniLM-L-6-v2444https://huggingface.co/microsoft/ms-marco-MiniLM-L-6-v2, fine-tuned on MS MARCO for passage ranking. We concatenate the IUR with the knowledge context, score the sequence using the model, and assign a World-Understanding rating of 10 if the the score exceeds 0.5 and 0 otherwise.
@@
While achieving high inter-annotator agreement (IAA) for subjective measures like World-Understanding and Unambiguity is inherently challenging, as evidenced by prior work showing human annotators struggling to exceed 30% IAA for related subjective criteria in NLG tasks karpinska-etal-2021-perils-2, we find that LLM-based proxy evaluation models, particularly GPT-3.5 and GPT-4, demonstrate considerable agreement with human raters for our task. Nonetheless, there remains scope for further boosting performance through additional prompt engineering and experimentation with adaptive strategies for selecting in-context examples. The prompts used for training both proprietary and open-source LLM proxy evaluator models are provided in Appendix B.
@@ 8 Related Work
The initiative to develop the IUR generation task springs from a need to reduce the brittleness of DST models. Cho et al. (2022) empirically demonstrate the brittleness of commonly used DST models by showing that their performance degrades in the face of various types of perturbations involving linguistic variations, coreferences, named entity references, paraphrases, and speech disfluencies. More generally, Zarcone et al. (2021) critique the academic community’s prevailing focus on incremental advancements on synthetic benchmarks for tasks such as DST, referred to as “playing the SNIPS game,” which often overlooks deeper issues regarding dataset realism.
@@
The process of generating IUR s bears resemblance to paraphrase generation (Zhou and Bhat, 2021) in the aspect of semantically preserving text transformation. IUR generation also shares an inverse conceptual similarity with the NLI task of inferring entailment from a premise and a hypothesis. In contrast, IUR generation can be thought of as generating an NLI premise given a hypothesis and a positive entailment class. Although Shen et al. (2018) explore this very task formulation, their work differs significantly from ours as it is not situated in a dialogue context.
@@ 9 Limitations and Future Work
Our proposed notion of IUR applies only to categorical slots with a small, fixed number of possible values (< 5), but not to those slots that can take on a large number of values. Future work can investigate the IUR generation task for such challenging dialogue schemas. We have also limited ourselves to supervised fine-tuning of LLMs. However, there is a rich literature on the use of reinforcement learning to guide language models towards specific text styles and content types, especially for abstract concepts of the likes of indirectness, which can be explored as future work Kaufmann et al. (2023). As Bowman and Dahl (2021) suggest, the ultimate evaluation measure for any NLP task should be grounded in in carefully annotated real user data. While modeling specific phenomena such as indirectness moves the needle on specific dialogue paradigms such as task-oriented dialogues, the community needs to evolve novel evaluation paradigms in the long run for wider forms of dialogue (Mannekote, 2023). Until then, works such as ours will continue to inform gaps in existing models.
@@ 10 Conclusion
As the research and development of task-oriented dialogue systems advances, there is a pressing need to bridge the gap between benchmark corpora and utterances “in the wild.” In our study, we concentrate on the phenomenon of “indirectness.” This occurs when a user conveys their desired outcome in a manner that requires the listener to utilize their general knowledge to deduce the intended value. We develop a multi-stage LLM-based pipeline to generate IndirectRequests, a dataset of IURs based on the schemas borrowed from the SGD dataset. IndirectRequests supplements existing benchmarks to evaluate NLU and DST models on realistic, indirect user requests lacking explicit slot values. Experiments with a state-of-the-art DST model validate the challenging nature of IndirectRequests. More broadly, our benchmark dataset can support future efforts for tasks such as API prediction, DST, NLU, which can lead to an overall improvement in the usability of virtual assistants for end users.

### v2

# Making Task-Oriented Dialogue Datasets More Natural by Synthetically Generating Indirect User Requests

@@ Abstract
Indirect User Requests (IURs), such as "It’s cold in here" instead of "Could you please increase the temperature?" are common in human-human task-oriented dialogue and require world knowledge and pragmatic reasoning from the listener. While large language models (LLMs) can handle these requests effectively, smaller models deployed on virtual assistants often struggle due to resource constraints. Moreover, existing task-oriented dialogue benchmarks lack sufficient examples of complex discourse phenomena such as indirectness. To address this, we propose a set of linguistic criteria along with an LLM-based pipeline for generating realistic IURs to test Natural Language Understanding (NLU) and Dialogue State Tracking (DST) models before deployment in a new domain. We also release IndirectRequests, a dataset of IURs based on the Schema Guided Dialog (SGD) corpus, as a comparative testbed for evaluating the performance of smaller models in handling indirect requests.
@@ 1 Introduction
Non-literal, indirect utterances are common in human-human task-oriented dialogue and require pragmatic understanding and world knowledge for successful interpretation (e.g., “It’s cold in here” instead of “Could you please increase the temperature?”) (Briggs and Scheutz, 2017). This phenomenon is a key area of interest in discourse pragmatics (Blum-Kulka and Hamo, 2011; Schegloff, 1999), supported by theoretical frameworks such as Grice’s maxims (Grice, 1975) and RST (Mann and Thompson, 1988). Figure 1 illustrates two instances of Indirect User Requests (IURs).
@@
Despite the prevalence of indirect utterances in everyday discourse and the human-level Natural Language Understanding (NLU) performance demonstrated by state-of-the-art large language models (LLMs) like GPT-4 (Achiam et al., 2023), current virtual assistants struggle to handle such utterances seamlessly (Mavrina et al., 2022). This can be attributed, in part, to the high computational cost associated with using state-of-the-art, large models for inference (Samsi et al., 2023; Sardana and Frankle, 2023). A common workaround is to employ smaller, cost-effective, task-specific models (Hsieh et al., 2023). However, this approach often compromises the generalizability and robustness offered by larger models.
Over the years, several benchmark datasets for task-oriented dialogue, such as MultiWOZ (Budzianowski et al., 2018), Schema Guided Dialog (SGD) (Rastogi et al., 2020), and FRAMES (Asri et al., 2017), have been curated by the dialogue systems community. However, these datasets have two key limitations that hinder their effectiveness in training smaller NLU models. First, their static nature and limited domain coverage make it difficult to evaluate NLU or Dialogue State Tracking (DST) models in new domains. Second, the controlled laboratory settings in which these datasets are crowdsourced lead to a distributional mismatch between the benchmark datasets and “in-the-wild” utterances Zarcone et al. (2021).
2 Schema-Guided Dialogue
To bridge this distributional gap, we present an LLM-based data generation pipeline to scalably generate IURs for a new task-oriented dialogue domain. Our work makes the following contributions:
@@ 2 Schema-Guided Dialogue
    We develop a set of linguistic criteria to formalize the concept of what constitutes an indirect user request in a task-oriented dialogue setting.
@@
    Finally, we empirically demonstrate the increased difficulty of the IURs by showing that the performance of a T5-based (Roberts et al., 2019) DST model significantly degrades when applied on IndirectRequests utterances as compared to their counterparts from SGD.
Before outlining the linguistic criteria, we first describe the paradigm of “schema-guided dialogue” since it serves as the basis for the task formulation.
@@ 4 The IndirectRequests Dataset
The goal of IUR generation is to take a domain, a domain schema (containing a user intent and a list of possible slot values), and a target slot value as inputs and output an IUR. The IUR, on its part, is expected to adhere to certain “linguistic criteria” to be valid.
@@ 4.3 Dataset Splits
Based on the crowdsourced labels for both Unambiguity and World-Understanding, we curate the IndirectRequests dataset and release it for public use. [^1] URL hidden for peer review. In going from the “raw” crowdsourced samples to the dataset, we split the dataset and systematically create labels for each sample for both Unambiguity and World-Understanding criteria. While splitting IndirectRequests into train, validation, and test sets, we split our samples based on same lines on which the services are split across the SGD dataset. This alignment with the SGD dataset splits is intended to aid future work that might need to compare our results with previous work reporting on SGD.
@@ 5.1 Proxy Evaluation Results
For the small LM category, we employ BERT-based models in a zero-shot setup. For the Unambiguity criterion, we frame the evaluation as $k$ Natural Language Inference (NLI) problems, where $k$ is the number of possible slot values. Each problem considers the candidate IUR as the premise and a possible slot value as the hypothesis. We use a BERT-based NLI model222nli-deberta-v3-small to obtain entailment scores and return the argmax score. If the maximum score is below 0.3, we deem the IUR ambiguous for that slot. For World-Understanding, we use ms-marco-MiniLM-L-6-v2333https://huggingface.co/microsoft/ms-marco-MiniLM-L-6-v2, fine-tuned on MS MARCO for passage ranking. We concatenate the IUR with the knowledge context, score the sequence using the model, and assign a World-Understanding rating of 10 if the the score exceeds 0.5 and 0 otherwise.
@@
While achieving high inter-annotator agreement (IAA) for subjective measures like World-Understanding and Unambiguity is inherently challenging, as evidenced by prior work showing human annotators struggling to exceed 30% IAA for related subjective criteria in NLG tasks Karpinska et al. (2021), we find that LLM-based proxy evaluation models, particularly GPT-3.5 and GPT-4, demonstrate considerable agreement with human raters for our task. Nonetheless, there remains scope for further boosting performance through additional prompt engineering and experimentation with adaptive strategies for selecting in-context examples. The prompts used for training both proprietary and open-source LLM proxy evaluator models are provided in Appendix B.
@@ 8 Related Work
The initiative to develop the IUR generation task springs from a need to reduce the brittleness of smaller NLU and DST models. Cho et al. (2022) empirically demonstrate the brittleness of commonly-used, small LM-based DST models by showing that their performance degrades in the face of various types of perturbations involving linguistic variations, coreferences, named entity references, paraphrases, and speech disfluencies. More generally, Zarcone et al. (2021) critique the academic community’s prevailing focus on incremental advancements on synthetic benchmarks for tasks such as DST, referred to as “playing the SNIPS game,” which often overlooks deeper issues regarding dataset realism.
@@
IUR generation is similar to paraphrase generation (Zhou and Bhat, 2021) in that both tasks are form of semantically-preserving text transformations. In fact, IUR generation can be viewed as the task of generating a highly specific form of paraphrase (that adheres to our three linguistic criteria). It can also be viewed as the inverse of the NLI task, where the objective is to generate a premise entailing a given hypothesis, rather than inferring entailment from a premise-hypothesis pair, albeit in a different context from Shen et al. (2018). Most closely related to our work, Ge et al. (2022) propose linguistic criteria based on Gricean Maxims (Grice, 1975) for the task of generating follow-up questions for interactive surveys. While both tasks prioritize relevance and coherence, they differ in their objectives: the former aims to elicit information from the user, while the latter focuses on clarity and unambiguity in conveying requests, often serving as the initial turn or an independent subdialogue thread.
@@ 9 Limitations and Future Work
We have limited ourselves to supervised fine-tuning of LLMs. However, there is a rich literature on the use of reinforcement learning to guide language models towards specific text styles and content types, especially for abstract concepts of the likes of indirectness, which can be explored as future work Kaufmann et al. (2023).
As Bowman and Dahl (2021) suggest, the ultimate evaluation measure for any NLP task should be grounded in in carefully annotated real user data. While modeling specific phenomena such as indirectness moves the needle on specific dialogue paradigms such as task-oriented dialogues, the community needs to evolve novel evaluation paradigms in the long run for wider forms of dialogue (Mannekote, 2023).
Finally, the linguistic criteria we have established for generating indirect requests in IndirectRequests are not only effective for the current dataset, but also serve as a robust and generalizable framework that can be leveraged in future work to create even more challenging and diverse datasets. For instance, by expanding the number of possible slot values per sample to tens or even hundreds, researchers can construct more complex and realistic datasets that push the boundaries of current NLU models.
@@ 10 Conclusion
In conclusion, our study addresses the gap between benchmark corpora and real-world utterances in task-oriented dialogue systems by focusing on the phenomenon of indirectness. We present a multi-stage LLM-based pipeline to generate IndirectRequests, a dataset of IURs based on the schemas from the SGD dataset. IndirectRequests complements existing benchmarks, enabling the evaluation of NLU and DST models on realistic, indirect user requests that lack explicit slot values. Experiments with a state-of-the-art DST model confirm the challenging nature of IndirectRequests. Furthermore, our data generation pipeline provides a versatile and efficient method for creating evaluation datasets for various task-oriented dialogue tasks on-the-fly, potentially driving significant improvements in the usability and performance of virtual assistants for the benefit of end users.